dy/dx= y(xy^(5) − 1)

1
Expert's answer
2022-03-22T18:40:11-0400

"y\u2019=y(xy^5-1)=xy^6-y"

"z=y^{-5}"

"y’=-\frac15z^{-\frac65}z’"

"-\frac15z^{-\frac65}z’=x z^{-\frac65}-z^{-\frac15}"

"z\u2019-5z=-5x"

"z=uv"

"u\u2019v+uv\u2019-5uv=-5x"

"u\u2019v+u(v\u2019-5v)=-5x"

"v\u2019-5v=0"

"v=e^{5x}"

"u\u2019e^{5x}=-5x"

"u=-\int 5xe^{-5x}dx=\frac15e^{-5x}(5x+1)+C"

"z=uv=x+\frac15+Ce^{5x}=y^{-5}"

"y=(x+\frac15+Ce^{5x})^{-\frac15}"

Answer: "y=(x+\frac15+Ce^{5x})^{-\frac15}".